OpenStudy (anonymous):

find dy/dx by implicit differentiation: siny=x^2-e^y

myininaya (myininaya):

\[\frac{d}{dx}(\sin(f(x))=?\] \[\frac{d}{dx}(x^2)=?\] \[\frac{d}{dx}e^{f(x)}=?\]

myininaya (myininaya):

The first and last problem I just listed, you will need to know chain rule. The middle one is just applying the power rule. The first one along with chain rule you will need to know the derivative of sin(x) is cos(x). The last one along with chain rule you will need to know the derivative of e^x is e^x. (f(x))'=f'(x) or df(x)/dx or df/dx or dy/dx

myininaya (myininaya):

Here are some of examples of chain rule: --- \[\frac{d}{dx}\cos(x^2+1)=(x^2+1)'(-\sin(x^2+1))=(2x+0)(-\sin(x^2+1))=\] \[=(2x)(-\sin(x^2+1))=-2xsin(x^2+1) \] --- \[\frac{d}{dx}\sin(x^6)=(x^6)'\cos(x^6)=6x^5 \cos(x^6)\] --- \[\frac{d}{dx} (f \circ g)(x)=\frac{d}{dx}f(g(x))=g'(x) \cdot f'(g(x)) \text{ or } =g'(x)(f' \circ g)(x)\]

myininaya (myininaya):

Do you think you can attempt answering the 3 problems I stated in my fist post?

OpenStudy (anonymous):

my main question is does e^y turn into e^x when using implicit differentiation? or does it change to e^y*y' and i have to solve from there?

myininaya (myininaya):

The derivative of y is y'. (e^y)'=y' * e^y

myininaya (myininaya):

or you can write it as e^y * y' like you have done so.

OpenStudy (anonymous):

Ok, so far what I have for the equation is: siny=x^2-e^y cosy*y'=2x-e^y*y' I'm stuck in solving the rest of the problem because of the e^y*y'. I don't know what the next step would be.

myininaya (myininaya):

Put your terms that have y' on one side of the equation and the other terms that don't contain y' on the on the side. Just like if we have 5x=5-8x We get our like terms together: 5x+8x=5 Factor x out like this: x(5+8)=5 Then divide both sides by 5+8 So we get x=5/(5+8)

OpenStudy (anonymous):

Ok I have: cosy*y'+e^y*y'=2x Now how do I get the e^y*y' as just y' so I can divide the cosy to get y' by itself.

myininaya (myininaya):

Did you look at my example any? If you are able to follow my example, you will see what to do.

myininaya (myininaya):

Like if not, like tell me what you don't get. Like I have 5x+8x=5 I want to solve for x. I factor the x out like: x(5+8)=5 Then to undo the multiplication of (5+8) by dividing both sides by (5+8).

OpenStudy (anonymous):

So the result will be: y'=2x/cosy+e^y

myininaya (myininaya):

y'=2x/(cosy+e^y) <---is correct.

OpenStudy (anonymous):

Ok, thank you so much for your help.

myininaya (myininaya):

Np. Do you understand what I was doing with my linear equation? Do you see how knowing how to solve a linear equation applies here?

OpenStudy (anonymous):

yea, I understood how you did all of that perfectly. Was just confused because e is raised to the y times y' power. I thought I would have to do something else to eliminate y' as part of the power for e.

myininaya (myininaya):

Oh no y' is not part of the exponent I thought you said (e^y)'=y'*e^y

myininaya (myininaya):

not e^(y*y') <--this is a big no no)

myininaya (myininaya):

\[\frac{d}{dx}e^{u(x)}=u'(x) \cdot e^{u(x)}\]

OpenStudy (anonymous):

oh ok. Thank you, the exponential stuff always confuses me.

myininaya (myininaya):

Well I hope it is less confusing now.

OpenStudy (anonymous):

i'm slowly getting it

myininaya (myininaya):

Let me know what questions you have.

OpenStudy (anonymous):

will do, thank you

OpenStudy (anonymous):

ok can you verify if I made a mistake somewhere: question is use implicit differentiation to find dy/dx: x^3+xy=y^2+x 3x^2+[x(y*y')+(1)(y)]=2y*y'+(1) 3x^2+xy*y'+y=2y*y'+1 3x^2+y-1=2y*y'-xy*y' 3x^2=y-1/2y-xy=(2y-xy)y'/2y-xy y'=3x2=y-1/2y-xy

myininaya (myininaya):

(xy)'=(x)'y+x(y)' =1 y+x y'

OpenStudy (anonymous):

oops i meant 3x^2+y-1/2y-xy

OpenStudy (anonymous):

is the solution i have down correct? (the second one i sent) or do i need to fix something?

myininaya (myininaya):

Well I think you still didn't apply what I said. You denominator is a little off because of it. And all I think you mean (3x^2+y-1)/(2y-xy) right?

OpenStudy (anonymous):

Yes that's what I meant.

myininaya (myininaya):

Did you look at what I said about the product rule? (xy)'=(x)' * y + x * (y)' Remember derivative or y is just y' not y*y' Just like if I said find y' for y=x+1, you would say y'=1 not y*y'=1 or y'=1/y The derivative of y is just y' (xy)' = (x)' * y + x * (y)' = 1 * y + x * y'

OpenStudy (anonymous):

oh ok, so fixing it i got: y'=(3x^2+y-1)/(2y-x)

myininaya (myininaya):

Beautiful.

OpenStudy (anonymous):

thank you. and I misread what you had typed, sorry. thank you so much for your help.

myininaya (myininaya):

About the product rule?

OpenStudy (anonymous):

yes

myininaya (myininaya):

Ok but after re-reading it, you understand the product rule and how you should apply if you have a product of functions. The only time you can ignore the product rule is if one of those functions or both are constant functions but yeah you could still use the product rule then. (xy)'=(x)' * y + x * (y)' .... assume c is constant like you know 5 and sqrt(3) and pi (these are all constants) 5 will always be 5. it doesn't change. that is why it is called constant, it is constantly 5 all the time. (cx)' = (c)' * x + c * (x)' = 0 *x + c* (x)' = c(x)' or just use constant multiple rule (cx)'=c(x)'

OpenStudy (anonymous):

Ok I have another one. find the derivative of the function: f(x)=4arcsin(x/2) I have so far: f'(x)=4[(1/2)/(squareroot 1-(x^2/4)) I don't know if I have this correct and stuck on what to do from here. Would by multiply the 4 and 1/2 to get: 2/(squareroot 1-(x^2/4))?

myininaya (myininaya):

\[f'(x)=4 \cdot \frac{\frac{1}{2}}{\sqrt{(1-(\frac{x}{2})^2})}=4 \cdot \frac{1}{2} \cdot \frac{1}{\sqrt{(1-\frac{x^2}{4}})}\] \[=2 \cdot \frac{1}{\sqrt{(\frac{4}{4}-\frac{x^2}{4})}}=\frac{2}{\sqrt{(\frac{4-x^2}{4})}}\] This should help. That is square root of[ (4-x^2)/4] For some reason the square root isn't displaying as I like.

myininaya (myininaya):

I didn't do the whole thing on simplifying. I left some of it to you. Let me know if you don't know what to do next.

OpenStudy (anonymous):

I don't know if this is right but I get \[\frac{ \sqrt{4+x^2} }{4-x^2 }\]

myininaya (myininaya):

How did you get that + sign? Also I think a multiple of 4 disappeared from your answer?

OpenStudy (anonymous):

I took the conjugate of \[\sqrt{4-x^2}\] to cancel the radical.

myininaya (myininaya):

\[\sqrt{4-x^2} \cdot \sqrt{4-x^2}=(4-x^2)=4-x^2\] just like \[\sqrt{9} \cdot \sqrt{9}=9 ;\sqrt{9} \cdot \sqrt{-9} \cancel{=} 9\] Also are you required to rationalize the bottom?

OpenStudy (anonymous):

ok as far as i can get then is 2/square root 4-x^2/2

OpenStudy (anonymous):

so would the answer be \[\frac{ 4\sqrt{4-x^2} }{ 4-x^2 }\]